2. Solve the system of equations shown below using the substitution method. Include your work. x - y = 5 1-3x + 3y = 2
Answer #1

x-y=5 - 3x + 3y = 2 2 x=sty 2 3 Put en een -3(5+y)+3y=2 -15-3y + 3y = 2 2 which is not true dience, no solution exists for giIf you have any questions do let me know in the comment section.

Please give a Thumbs up. Thanks in advance.

Stay Safe.